OpenStudy (anonymous):

Find the number halfway between 1/2 and 5/6

OpenStudy (anonymous):

A. 13/20 B. 27/40 C. 2/3 D. 7/10

OpenStudy (rock_mit182):

so you hade to add them up and then, the result diveded by two.

OpenStudy (anonymous):

So is it C? Because I turned them all into decimals.

OpenStudy (rock_mit182):

why don't you get the commun factor on the denominator and add them up, i mean they are working with fractions

OpenStudy (anonymous):

Ok... :P

OpenStudy (rock_mit182):

;)

OpenStudy (anonymous):

I added them together and got 1 2/6

Miracrown (miracrown):

Then we need to divide that by 2

OpenStudy (anonymous):

How do I do that again?

Miracrown (miracrown):

that is how we get the midpoint we add the 2 numbers then divide by 2

Miracrown (miracrown):

\[\frac{ 1 }{ 2 } + \frac{ 5 }{ 6 } = \frac{ 3 }{ 6 } + \frac{ 5 }{ 6 } = \frac{ 8 }{ 6 }\]

OpenStudy (rock_mit182):

yeah thats correct

Miracrown (miracrown):

How do you add them, or how do you divide by 2? Are you asking about the dividing by 2 part? You found the 1 2/6 from adding 1/2 and 5/6 How did you find that?

OpenStudy (anonymous):

I'm asking about the dividing, I understand the adding.

Miracrown (miracrown):

Did you find the common denominator? I just put the addition of 1/2 and 5/6

OpenStudy (rock_mit182):

you have to check ypur adding ;) in fraction

OpenStudy (anonymous):

Yeah.. 6.

Miracrown (miracrown):

Ok... We can leave the 1 2/6 as a fraction to make the dividing easier so 8/6

OpenStudy (anonymous):

can someone help @made pleaseeeeeeeeeeee :p

Miracrown (miracrown):

Then to divide 8/6 by 2, we multiply by 1/2

OpenStudy (rock_mit182):

oh i see well you did it right is jjust another way to expressed that

OpenStudy (anonymous):

I will give medal

Miracrown (miracrown):

Then to divide 8/6 by 2, we multiply by 1/2 Do you follow that part? 8/6 divided by 2?

OpenStudy (anonymous):

I got the answer. :)

OpenStudy (rock_mit182):

awesome !

OpenStudy (anonymous):

C. :)

Miracrown (miracrown):

great! :]

Miracrown (miracrown):

Yes, 2/3 is the correct answer